Question-4


(a)  Discuss how potential difference and electric field strength are related. Give an example


(b)  What is the strength of the electric field in a region where the electric potential is constant?   


(c)   What is the relationship between voltage and energy? More precisely, what is the relationship between potential difference and electric potential energy?


1
Expert's answer
2021-01-16T17:22:29-0500

a) In equation form, the general relationship between voltage and electric field is


"E=-\frac{\Delta V}{\Delta s}"

where Δs is the distance over which the change in potential, ΔV, takes place. The minus sign tells us that E points in the direction of decreasing potential. The electric field is said to be the gradient (as in grade or slope) of the electric potential.


The potential difference or voltage between the plates of capacitor is


"V_{AB}=Ed=(3\cdot10^6)(0.025)=75\ kV"

b) The strength of the electric field in a region where the electric potential is constant is zero:


"E=-\frac{0}{\Delta s}=0\frac{V}{m}"


c)


"U_{AB}=qV_{AB}"